Calculate the change the entropy, Change in Gibbs free energy, change in Enthalpy and change in Volume if 0.5mol of Oxygen is mixed with 0.5mol of Nitrogen at 25 degree Celsius and 1 atm assuming they are Ideal gases.


1
Expert's answer
2021-04-13T02:08:44-0400

"N_2 +2O_2-->2NO_2"


Total mol = 0.5+ 0.5 = 1.0 mol

R = gas constant = 0.0821 atm•L/mol•K

Pressure = 1 atm


"Entropy = - 0.0821[ \frac{0.5}{1.0}×ln(\frac{0.5}{1.0})+ \frac{0.5}{1.0}×ln(\frac{0.5}{1.0})]"


"Entropy =-0.0821[-0.347+-0.347]"



"Entropy =-0.0821[-0.694] =0.057 j"


"Enthalpy = 5.7\u00d710^{-5} kJ"